A transverse sinusoidal wave is generated at one end of a long horizontal string by a bar that moves the end up and down through a distance of 2.6 cm. The motion is repeated regularly 150 times per second. a) If the distance between adjacent wave troughs is 16.0 cm, find the amplitude, frequency, speed and wavelength of the wave.


1
Expert's answer
2021-07-19T09:52:29-0400
"A=0.5(2.6)=1.3\ cm\\f=125\ Hz\\v=\frac{2\pi(125)}{0.16}=4910\frac{m}{s}
\\\lambda=16\ cm"